How many moles of Ag can be formed when 5.0 mol of Zn react with 4.2 mol of AgNO3?
1
Expert's answer
2013-02-13T09:09:56-0500
Firstly, let's find the amount of substance for each product:
n(AgNO3)= 4.2 mol
n(Zn) = 5 mol
Next reaction:
2AgNO3 + Zn = Zn(NO3)2 + 2Ag
zinc is in excess, so we'll use silver nitrate for calculations:
n(Ag) = n(AgNO3)
m(Ag) = n(Ag)*Ar(Ag)
m(Ag) = 4.2 mol * 108 g/mol
m(Ag) = 453,6 g